• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 15:29:36 on April28,2008

    Entry number 235533

    keyword=day shift summary

    Day shift summary:

    shift workers: J. P. Chen (SL), D. Armstrong (TO), X. Qian (3rd).

    8:00 production at Kin. L: LHRS at 12.451 degree, 1.007 GeV/c
    BigBite 105 degree, 390 A,
    At the beginning of shift, 0.144 C.
    11:00 RF problem. beam off. (run 4260).
    11:20 Beam back.
    12:55 Nikos started to turn on FPP HVs for FPP test.
    13:15 FPP HVs all set to 1750 V. Started FPP test (run 4273-4282).
    FPP HV settings: 1750, 1800, 1850, 1875, 1700, 1775, 1825, 1900V.
    14:17, MCC called that beam off to change Wien angle for Hall C
    (run 4279).
    14:30 beam back, Wien angles changed from -40 to -30 degree.
    14:53-14:56 Wien cangle change to -50 degree.
    14:56 turned off FPP HVs.
    15:20 MCC cchange Wien angle to -40 degree.
    Moller starts after the Wien angle change.

    run 4241/4242/4243/4244/4245/4246/4247/4248/4249/4250/4251
    4252/4253/4254/4255/4256/42574258/4259, good runs, 1M.
    run 4260 RF problem. 0.94M.
    run 4261/4262/4263/4264/4265/4266/4267/4268/4269/4270/4271/4272. 1M.

    Following runs also for FPP test, 1M events each run:
    run#, FPP HV
    4273, 1750
    4274, 1800
    4275, 1850
    4276, 1875
    4277, 1700
    4278, 1775
    4279, 1825, 252K, stopped for wien angle change from -40 degree to -30 degree (next run will be -30 degree).
    4280, 1825, wien angle -30 degree.
    4281, 1900.
    4282, 1900. 514K. stopped for another Wien angle change.
    FPP test ends.

    4283/4284, Wien angle -50 degree. 1M.
    4285. Wien -50 degree, 435K, stopped to change Wien angel back to -40 degree.